OpenStudy (antiprime):

f 9999 = 4, 8888 = 8, 1816 = 6, 1212 = 0, then 1919 =?

OpenStudy (antiprime):

after this i have to go

OpenStudy (antiprime):

ik the correct anwser

OpenStudy (error1603):

9999=4 9=1 8888=8 8=2 1816 = 3 x+2+x+y = 3 1212=0 1=0, 2=0 1816 = 3 6=2 1919 = 0+1+0+1 = 2

OpenStudy (antiprime):

no ur wrong

OpenStudy (antiprime):

its 4

OpenStudy (error1603):

Really got me

OpenStudy (antiprime):

Look at the numbers with closed areas, e.g the top of 9. The number 9999 has 4 closed areas. 8888 has 8 closed areas, the top and bottom parts of the 8. 1816 has 3 closed areas, (top and bottom of 8 and bottom of 6, and it has 2 other digits ( 3*2=6). 1212 has 0 closed areas (0*4=0). this is y

OpenStudy (error1603):

Oh I see

OpenStudy (antiprime):

yea i gtg byee

OpenStudy (error1603):

Bye